# Quantum Wave Simulation with Sources and Loss Functions

### v1

@@ Appendix C Finite Difference Discretization for the Acoustic Wave equation
In this Appendix section, we discuss the discretization of the wave equation using the FD method on a uniform, rectangular, and staggered grid [79]. We show how to build a pair of discrete gradient and divergence operators to ensure that $\hat{\mathbf{A}}^{T}=-\hat{\mathbf{A}}$ is anti-Hermitian (see Section 2). We also discuss the implementation of Dirichlet and Neumann BCs on arbitrary boundaries of the domain by employing the constraints reduction introduced in Appendix D.
Recall the acoustic wave equation in pressure-velocity formulation introduced in Section 2.1
$\hat{\mathbf{B}}_{\text{acoustic}}\frac{\partial\hat{\mathbf{w}}_{\text{acoustic}}}{\partial t}$ $=\hat{\mathbf{A}}_{\text{acoustic}}\hat{\mathbf{w}}_{\text{acoustic}}$
@@
$(\nabla u(\mathbf{x},t))_{i}$ $=\frac{\partial u}{\partial x_{i}}(\mathbf{x},t)\approx\frac{u(\mathbf{x}+\frac{\Delta x_{i}}{2}\mathbf{e}_{i},t)-u(\mathbf{x}-\frac{\Delta x_{i}}{2}\mathbf{e}_{i},t)}{\Delta x_{i}},\forall i\in D,$ (81)
$\nabla\cdot\mathbf{v}(\mathbf{x},t)$ $=\sum_{i\in D}\frac{\partial v_{i}}{\partial x_{i}}(\mathbf{x},t)\approx\sum_{i\in D}\frac{v_{i}(\mathbf{x}+\frac{\Delta x_{i}}{2}\mathbf{e}_{i},t)-v_{i}(\mathbf{x}-\frac{\Delta x_{i}}{2}\mathbf{e}_{i},t)}{\Delta x_{i}}.$ (82)
We now need to replace the continuous fields $u(\mathbf{x},t)\colon\mathbb{R}^{D}\times[0,T]\to\mathbb{R},\mathbf{v}(\mathbf{x},t)\colon\mathbb{R}^{D}\times[0,T]$ with their discretized counterparts $\mathbf{u}(t)\colon[0,T]\to\mathbb{R}^{N_{u}},\mathbf{v}(t)\colon[0,T]\to\mathbb{R}^{N_{v}}$ by employing a staggered grid where the velocities are staggered in between pressure points. We introduce the set of numbers of grid points $\mathcal{N}=\{N_{i}\mid i\in\{1,\ldots,D\},N_{i}=\text{\# of grid points in the $i^{\text{th}}$ direction}\}$. Since the velocity is a vector field, we need to distinguish between velocities in different dimensions: we will use $\mathbf{v}_{i}$ from now on to denote the velocity acting in the $i^{\text{th}}$ direction, such that $\mathbf{v}_{i}(t)\colon[0,T]\to\mathbb{R}^{N_{v_{i}}}$ where $N_{v_{i}}=(N_{i}-1)\prod_{j\neq i}N_{j}$. Notice that this discretization imposes the discretized velocity field in the $i^{\text{th}}$ direction to have one less Degree of Freedom (DOF) in that direction. The pressure field $\mathbf{u}$ is going to be discretized in $N_{u}=\prod_{i}N_{i}$ DOF and the full velocity field $\mathbf{v}=[\mathbf{v}_{1},\ldots,\mathbf{v}_{D}]$ in $N_{v}=\sum_{i}((N_{i}-1)\prod_{j\neq i}N_{j})$ DOF.
@@
$\mathcal{X}_{u}=\{(x,y)\in\Omega\mid x=x_{0}+(i-1)\Delta x,y=y_{0}+(j-1)\Delta y,\forall i\in\{1,\ldots,N_{x}\},\forall j\in\{1,\ldots,N_{y}\}\},$
@@
$\mathcal{X}_{v_{x}}=\{(x+\Delta x/2,y)\in\Omega\mid(x,y)\in\mathcal{X}_{u}\},$
@@
$\mathcal{X}_{v_{y}}=\{(x,y+\Delta y/2)\in\Omega\mid(x,y)\in\mathcal{X}_{u}\}.$
@@
$[\mathbf{u}]_{\ell}(t)$ $=u(\Phi_{u}(\ell),t),\forall\ell\in\{1,\ldots,N_{u}\},$
$[\mathbf{v}_{x}]_{k}(t)$ $=v_{x}(\Phi_{v_{x}}(k),t),\forall k\in\{1,\ldots,N_{v_{x}}\},$
$[\mathbf{v}_{y}]_{k-N_{v_{x}}}(t)$ $=v_{y}(\Phi_{v_{y}}(k),t),\forall k\in\{N_{v_{x}}+1,\ldots,N_{v_{x}}+N_{v_{y}}\}.$
@@ C.2 Discrete Gradient and Divergence Operators, and their Anti-Transpose Relationship
$(\nabla u)_{x}(\Phi_{v_{x}}(k))=\frac{\partial u}{\partial x}(\Phi_{v_{x}}(k))$ $\approx\frac{u(\Phi_{v_{x}}(k)+\frac{\Delta x}{2}\mathbf{e}_{x})-u(\Phi_{v_{x}}(k)-\frac{\Delta x}{2}\mathbf{e}_{x})}{\Delta x}$
@@
$(\nabla u)_{y}(\Phi_{v_{y}}(k))=\frac{\partial u}{\partial y}(\Phi_{v_{y}}(k))$ $\approx\frac{u(\Phi_{v_{y}}(k)+\frac{\Delta y}{2}\mathbf{e}_{y})-u(\Phi_{v_{y}}(k)-\frac{\Delta y}{2}\mathbf{e}_{y})}{\Delta y}$
@@
$\nabla\to\mathbf{G}=(\mathbf{G})_{k,\ell}=\begin{cases}\pm 1/\Delta x&\text{if}k\in\{1,\ldots,N_{v_{x}}\}\text{ and }\Phi_{u}(\ell)=\Phi_{v_{x}}(k)\pm\frac{\Delta x}{2}\mathbf{e}_{x},\\ \pm 1/\Delta y&\text{if }k\in\{N_{v_{x}}+1,\ldots,N_{v_{x}}+N_{v_{y}}\}\text{ and }\Phi_{u}(\ell)=\Phi_{v_{y}}(k)\pm\frac{\Delta y}{2}\mathbf{e}_{y},\\ 0&\text{otherwise},\end{cases}$ (83)
@@
$\nabla\cdot\mathbf{v}(\Phi_{u}(\ell))$ $=\frac{\partial v_{x}}{\partial x}(\Phi_{u}(\ell))+\frac{\partial v_{y}}{\partial y}(\Phi_{u}(\ell))$
$\approx\frac{v_{x}(\Phi_{u}(\ell)+\frac{\Delta x}{2}\mathbf{e}_{x})-v_{x}(\Phi_{u}(\ell)-\frac{\Delta x}{2}\mathbf{e}_{x})}{\Delta x}+\frac{v_{y}(\Phi_{u}(\ell)+\frac{\Delta y}{2}\mathbf{e}_{y})-v_{y}(\Phi_{u}(\ell)-\frac{\Delta y}{2}\mathbf{e}_{y})}{\Delta y}$
$\approx\frac{v_{x}(\Phi_{v_{x}}(k_{x+}))-v_{x}(\Phi_{v_{x}}(k_{x-}))}{\Delta x}+\frac{v_{y}(\Phi_{v_{y}}(k_{y+}))-v_{y}(\Phi_{v_{y}}(k_{y-}))}{\Delta y}$
$\approx\frac{[\mathbf{v}_{x}]_{k_{x+}}-[\mathbf{v}_{x}]_{k_{x-}}}{\Delta x}+\frac{[\mathbf{v}_{y}]_{k_{y+}}-[\mathbf{v}_{y}]_{k_{y-}}}{\Delta y},$
@@
$\nabla\cdot\to\mathbf{D}=(\mathbf{D})_{\ell,k}=\begin{cases}\pm 1/\Delta x&\text{if }k\in\{1,\ldots,N_{v_{x}}\}\text{ and }\Phi_{v_{x}}(k)=\Phi_{u}(\ell)\pm\frac{\Delta x}{2}\mathbf{e}_{x},\\ \pm 1/\Delta y&\text{if }k\in\{N_{v_{x}}+1,\ldots,N_{v_{x}}+N_{v_{y}}\}\text{ and }\Phi_{v_{y}}(k)=\Phi_{u}(\ell)\pm\frac{\Delta y}{2}\mathbf{e}_{y},\\ 0&\text{otherwise}.\end{cases}$ (84)
The anti-transpose property $\mathbf{G}=-\mathbf{D}^{T}$ can be verified by showing that $(\mathbf{G})_{k,\ell}=-(\mathbf{D})_{\ell,k}$ for all $\ell,k$. This is easily checked by observing that the only difference in the definition of the operators $\mathbf{G}$ and $\mathbf{D}$ is the position of the mappings $\Phi_{u}$, $\Phi_{v_{x}}$ (respectively, $\Phi_{u}$, $\Phi_{v_{y}}$) in the conditions $\Phi_{u}(\ell)=\Phi_{v_{x}}(k)\pm\frac{\Delta x}{2}\mathbf{e}_{x}$ and $\Phi_{v_{x}}(k)=\Phi_{u}(\ell)\pm\frac{\Delta x}{2}\mathbf{e}_{x}$ (respectively, the conditions $\Phi_{u}(\ell)=\Phi_{v_{y}}(k)\pm\frac{\Delta y}{2}\mathbf{e}_{y}$ and $\Phi_{v_{y}}(k)=\Phi_{u}(\ell)\pm\frac{\Delta y}{2}\mathbf{e}_{y}$), which is the cause of sign flip of the components. This shows that the FD discretization using a uniform, rectangular, staggered grid produces a pair of gradient and divergence operators leading to an anti-Hermitian matrix $A=\begin{bmatrix}\mathbf{0}_{N_{u}\times N_{u}}&-D\\ -G&\mathbf{0}_{D\times D}\end{bmatrix}$.
@@ C.3 Implementation of Boundary Conditions
Since these added constraints result in a system where $\mathbf{R}_{f}=0$ (see Appendix D), i.e., the unconstrained DOF do not appear in the constraint matrix $\mathbf{R}$, the modified matrices will be $\underline{\mathbf{B}}=\mathbf{B}_{ff}$ and $\underline{\mathbf{A}}=\mathbf{A}_{ff}$, which are the original with some rows and columns removed corresponding to the newly constrained DOF.
@@ Appendix D Generalizing Boundary Conditions to Linear Constraints
where $\mathbf{R}\in\mathbb{R}^{N_{c}\times N}$ is the constraint matrix, $N_{c}$ is the number of constraints, $\mathbf{w}_{f}\in\mathbb{R}^{N-N_{c}}$ represents the free (unconstrained) DOF, and $\mathbf{w}_{c}\in\mathbb{R}^{N_{c}}$ represents the constrained DOF, which will be eliminated from the equations of motion. The vector $\mathbf{b}(t)$ is non-zero if the constraints are non-homogeneous. For example, in a 2D acoustic wave problem with FD discretization, homogeneous Dirichlet (Neumann) BCs along any boundary orthogonal to the $x$ or $y$ velocity components can be enforced by setting $\mathbf{b}_{i}=0$, $\mathbf{R}_{f}=\mathbf{0}$ and $\mathbf{R}_{c}=\mathds{1}$ for all pressure (velocity) values on the boundary.
To incorporate the BCs (or constraints) into the wave equation, we follow the approach in [48]. We solve for the constrained DOF
@@
Rearranging the equations in this way preserves the Hermitian or anti-Hermitian nature of the matrices. Substituting (87) into the above equation and solving for the free DOF yields
@@
In general, this source cannot be implemented within our framework as the resulting wave field that needs to be initialized is neither radially symmetric nor populates a volume that is independent of $N$. However, if $\mathbf{b}(t)$ is pulse-like and a point source in space, then the logic of Section 4.1 applies, giving rise to a quantum initialization that is independent on $N$. Hence, certain boundary or displacement control problems can be implemented within the presented framework.

### v3

@@ Appendix C Finite Difference Discretization for the Acoustic Wave equation
As an example, in this Appendix, we discuss the discretization of the acoustic wave equation using the FD method on a uniform, rectangular, and staggered grid [82]. We show how to build a pair of discrete gradient and divergence operators to ensure that $\hat{\mathbf{A}}^{T}=-\hat{\mathbf{A}}$ is anti-Hermitian (see Section II). We also discuss the implementation of Dirichlet and Neumann BCs on arbitrary boundaries of the domain by employing the constraints reduction introduced in Appendix D.
Recall the acoustic wave equation in pressure-velocity formulation introduced in Section II.1
$\hat{\mathbf{B}}_{\text{acoustic}}\frac{\partial\hat{\mathbf{w}}_{\text{acoustic}}}{\partial t}$ $=\hat{\mathbf{A}}_{\text{acoustic}}\hat{\mathbf{w}}_{\text{acoustic}}$ (72)
@@
$(\nabla u(\mathbf{x},t))_{i}$ $=\frac{\partial u}{\partial x_{i}}(\mathbf{x},t)$ (74)
$\approx\frac{u(\mathbf{x}+\frac{\Delta x_{i}}{2}\mathbf{e}_{i},t)-u(\mathbf{x}-\frac{\Delta x_{i}}{2}\mathbf{e}_{i},t)}{\Delta x_{i}},$
and
$\nabla\cdot\mathbf{v}(\mathbf{x},t)$ $=\sum_{i\in D}\frac{\partial v_{i}}{\partial x_{i}}(\mathbf{x},t)$ (75)
$\approx\sum_{i\in D}\frac{v_{i}(\mathbf{x}+\frac{\Delta x_{i}}{2}\mathbf{e}_{i},t)-v_{i}(\mathbf{x}-\frac{\Delta x_{i}}{2}\mathbf{e}_{i},t)}{\Delta x_{i}}.$
We now need to replace the continuous fields $u(\mathbf{x},t)\colon\mathbb{R}^{D}\times[0,T]\to\mathbb{R},\mathbf{v}(\mathbf{x},t)\colon\mathbb{R}^{D}\times[0,T]$ with their discretized counterparts $\mathbf{u}(t)\colon[0,T]\to\mathbb{R}^{N_{u}},\mathbf{v}(t)\colon[0,T]\to\mathbb{R}^{N_{v}}$ by employing a staggered grid where the velocities are staggered in between pressure points. We introduce the set of numbers of grid points $\mathcal{N}=\{N_{i}\mid i\in\{1,\ldots,D\},N_{i}=\text{\# of grid points in the $i^{\text{th}}$ direction}\}$. Since the velocity is a vector field, we need to distinguish between velocities in different dimensions: we will use $\mathbf{v}_{i}$ from now on to denote the velocity acting in the $i^{\text{th}}$ direction, such that $\mathbf{v}_{i}(t)\colon[0,T]\to\mathbb{R}^{N_{v_{i}}}$ where $N_{v_{i}}=(N_{i}-1)\prod_{j\neq i}N_{j}$. Notice that this discretization requires the discretized velocity field in the $i^{\text{th}}$ direction to have one less DOF in that direction. The pressure field $\mathbf{u}$ is discretized in $N_{u}=\prod_{i}N_{i}$ DOFs and the full velocity field $\mathbf{v}=[\mathbf{v}_{1},\ldots,\mathbf{v}_{D}]$ in $N_{v}=\sum_{i}((N_{i}-1)\prod_{j\neq i}N_{j})$ DOFs.
@@
$\mathcal{X}_{u}=$$\{(x,y)\in\Omega\mid$ (76)
$x=x_{0}+(i-1)\Delta x,\quad y=y_{0}+(j-1)\Delta y,$
$\forall i\in\{1,\ldots,N_{x}\},\quad\forall j\in\{1,\ldots,N_{y}\}\},$
@@
$\mathcal{X}_{v_{x}}=\{(x+\Delta x/2,y)\in\Omega\mid(x,y)\in\mathcal{X}_{u}\},$ (77)
@@
$\mathcal{X}_{v_{y}}=\{(x,y+\Delta y/2)\in\Omega\mid(x,y)\in\mathcal{X}_{u}\}.$ (78)
@@
${}_{\ell}(t)=$ $u(\Phi_{u}(\ell),t),$ (79)
$\forall\ell\in\{1,\ldots,N_{u}\},$
$[\mathbf{v}_{x}]_{k}(t)=$ $v_{x}(\Phi_{v_{x}}(k),t),$
$\forall k\in\{1,\ldots,N_{v_{x}}\},$
$[\mathbf{v}_{y}]_{k-N_{v_{x}}}(t)=$ $v_{y}(\Phi_{v_{y}}(k),t),$
$\forall k\in\{N_{v_{x}}+1,\ldots,N_{v_{x}}+N_{v_{y}}\}.$
@@ C.2 Discrete Gradient and Divergence Operators, and their Anti-Transpose Relationship
$(\nabla u)_{x}(\Phi_{v_{x}}(k))$ $=\frac{\partial u}{\partial x}(\Phi_{v_{x}}(k))$ (80)
$\approx\frac{u(\Phi_{v_{x}}(k)+\frac{\Delta x}{2}\mathbf{e}_{x})}{\Delta x}$
$-\frac{u(\Phi_{v_{x}}(k)-\frac{\Delta x}{2}\mathbf{e}_{x})}{\Delta x}$
@@
$(\nabla u)_{y}(\Phi_{v_{y}}(k))$ $=\frac{\partial u}{\partial y}(\Phi_{v_{y}}(k))$ (81)
$\approx\frac{u(\Phi_{v_{y}}(k)+\frac{\Delta y}{2}\mathbf{e}_{y})}{\Delta y}$
$-\frac{u(\Phi_{v_{y}}(k)-\frac{\Delta y}{2}\mathbf{e}_{y})}{\Delta y}$
@@
$\nabla\to\mathbf{G}$ $=(\mathbf{G})_{k,\ell}$ (82)
={±$=\begin{cases}\pm 1/\Delta x&\begin{aligned} &\text{if }k\in\{1,\ldots,N_{v_{x}}\}\\ &\text{and }\Phi_{u}(\ell)=\Phi_{v_{x}}(k)\pm\frac{\Delta x}{2}\mathbf{e}_{x},\end{aligned}\\ \pm 1/\Delta y&\begin{aligned} &\text{if }k\in\{N_{v_{x}}+1,\ldots,N_{v_{x}}+N_{v_{y}}\}\\ &\text{and }\Phi_{u}(\ell)=\Phi_{v_{y}}(k)\pm\frac{\Delta y}{2}\mathbf{e}_{y},\end{aligned}\\ 0&\text{otherwise},\end{cases}$
@@
$\nabla\cdot\mathbf{v}(\Phi_{u}(\ell))$ $=\frac{\partial v_{x}}{\partial x}(\Phi_{u}(\ell))+\frac{\partial v_{y}}{\partial y}(\Phi_{u}(\ell))$ (83)
$\approx\frac{v_{x}(\Phi_{u}(\ell)+\frac{\Delta x}{2}\mathbf{e}_{x})-v_{x}(\Phi_{u}(\ell)-\frac{\Delta x}{2}\mathbf{e}_{x})}{\Delta x}$
$+\frac{v_{y}(\Phi_{u}(\ell)+\frac{\Delta y}{2}\mathbf{e}_{y})-v_{y}(\Phi_{u}(\ell)-\frac{\Delta y}{2}\mathbf{e}_{y})}{\Delta y}$
$\approx\frac{v_{x}(\Phi_{v_{x}}(k_{x+}))-v_{x}(\Phi_{v_{x}}(k_{x-}))}{\Delta x}$
$+\frac{v_{y}(\Phi_{v_{y}}(k_{y+}))-v_{y}(\Phi_{v_{y}}(k_{y-}))}{\Delta y}$
$\approx\frac{[\mathbf{v}_{x}]_{k_{x+}}-[\mathbf{v}_{x}]_{k_{x-}}}{\Delta x}$
$+\frac{[\mathbf{v}_{y}]_{k_{y+}}-[\mathbf{v}_{y}]_{k_{y-}}}{\Delta y},$
@@
$\nabla\cdot\to\mathbf{D}$ $=(\mathbf{D})_{\ell,k}$ (84)
={±$=\begin{cases}\pm 1/\Delta x&\begin{aligned} &\text{if }k\in\{1,\ldots,N_{v_{x}}\}\\ &\text{and }\Phi_{v_{x}}(k)=\Phi_{u}(\ell)\pm\frac{\Delta x}{2}\mathbf{e}_{x},\end{aligned}\\ \pm 1/\Delta y&\begin{aligned} &\text{if }k\in\{N_{v_{x}}+1,\ldots,N_{v_{x}}+N_{v_{y}}\}\\ &\text{and }\Phi_{v_{y}}(k)=\Phi_{u}(\ell)\pm\frac{\Delta y}{2}\mathbf{e}_{y},\end{aligned}\\ 0&\text{otherwise}.\end{cases}$
The anti-transpose property $\mathbf{G}=-\mathbf{D}^{T}$ can be verified by showing that $(\mathbf{G})_{k,\ell}=-(\mathbf{D})_{\ell,k}$ for all $\ell,k$. This is easily checked by observing that the only difference in the definition of the operators $\mathbf{G}$ and $\mathbf{D}$ is the position of the mappings $\Phi_{u}$, $\Phi_{v_{x}}$ (respectively, $\Phi_{u}$, $\Phi_{v_{y}}$) in the conditions $\Phi_{u}(\ell)=\Phi_{v_{x}}(k)\pm\frac{\Delta x}{2}\mathbf{e}_{x}$ and $\Phi_{v_{x}}(k)=\Phi_{u}(\ell)\pm\frac{\Delta x}{2}\mathbf{e}_{x}$ (respectively, the conditions $\Phi_{u}(\ell)=\Phi_{v_{y}}(k)\pm\frac{\Delta y}{2}\mathbf{e}_{y}$ and $\Phi_{v_{y}}(k)=\Phi_{u}(\ell)\pm\frac{\Delta y}{2}\mathbf{e}_{y}$), which is the cause of sign flip of the components. This shows that the FD discretization using a uniform, rectangular, staggered grid produces a pair of gradient and divergence operators leading to an anti-Hermitian matrix $\mathbf{A}=\begin{bmatrix}\mathbf{0}_{N_{u}\times N_{u}}&-\mathbf{D}\\ -\mathbf{G}&\mathbf{0}_{D\times D}\end{bmatrix}$.
@@ C.3 Implementation of Boundary Conditions
Since these added constraints result in a system where $\mathbf{R}_{f}=0$ (see Appendix D), i.e., the unconstrained DOFs do not appear in the constraint matrix $\mathbf{R}$, the modified matrices will be $\underline{\mathbf{B}}=\mathbf{B}_{ff}$ and $\underline{\mathbf{A}}=\mathbf{A}_{ff}$, which are the original with some rows and columns removed corresponding to the newly constrained DOFs.
@@ Appendix D Generalizing Boundary Conditions to Linear Constraints
where $\mathbf{R}\in\mathbb{R}^{N_{c}\times N}$ is the constraint matrix, $N_{c}$ is the number of constraints, $\mathbf{w}_{f}\in\mathbb{R}^{N-N_{c}}$ represents the free (unconstrained) DOFs, and $\mathbf{w}_{c}\in\mathbb{R}^{N_{c}}$ represents the constrained DOFs, which will be eliminated from the equations of motion. The vector $\mathbf{b}(t)$ is non-zero if the constraints are non-homogeneous. For example, in a 2D acoustic wave problem with FD discretization, homogeneous Dirichlet (Neumann) BCs along any boundary orthogonal to the $x$ or $y$ velocity components can be enforced by setting $\mathbf{b}_{i}=0$, $\mathbf{R}_{f}=\mathbf{0}$ and $\mathbf{R}_{c}=\mathds{1}$ for all pressure (velocity) values on the boundary.
To incorporate the BCs (or constraints) into the wave equation, we follow the approach in [52]. We solve for the constrained DOFs
@@
Rearranging the equations in this way preserves the Hermitian or anti-Hermitian nature of the matrices. Substituting Eq. (87) into the above equation and solving for the free DOFs yields
@@
In general, this source cannot be implemented within our framework as the resulting wave field that needs to be initialized is neither radially symmetric nor populates a volume that is independent of $N$. However, if $\mathbf{b}(t)$ is pulse-like and a point source in space, then the logic of Section V.1 applies, giving rise to a quantum initialization that is independent on $N$. Hence, certain boundary or displacement control problems can be implemented within the presented framework.